If 0.5icap+o.8jcap+ckcap is a unit vector.Find the value of C.

Grade:11

7 Answers

akash solanki
11 years ago

DEAR NIMISH

 

 If

0.5icap+o.8jcap+ckcap is a unit vector then its magnitude should be 1. So the value of C is 0.33


goutam kunal
11 years ago

111/2

check it it is 100% correct.

rupali sehgal
11 years ago

Hello Nimish

modulus of 0.5i+0.8j+C k= √(0.5)2 + (0.8)2 + C

                                   = √0.25+0.64+C2

                                   = √0.89 + C2

But its a unit vector. therefore it''s magnitude should be equal to 1.

√0.89+C2 = 1

squaring both sides,

0.89+C2= 1

C2= 1-0.89 = 0.11

C = 0.33

 

All the best!

Nimish Singh
11 years ago

Hey the answer was right but tell the way to solve it.

Abhishekh kumar sharma
10 years ago

0.25+.64+c^2=1

then c=(11)^1\2

Don
6 years ago
The formula is _/a^2+b^2+c^2 = Magnitude Put value in the above formula _/(0.5)^2+(0.8)^2+c^2=1By calculating we get -/0.11
ankit singh
3 years ago
|a+b+c|=1
√0.5²+0.8²+c²=1
squaring both sides 
0.5²+0.8²+c²=1
0.25+0.64+c²=1
c²+0.89=1.00
c²=0.11
c=√0.11